# Bayesian Inference for PDE-based Inverse Problems using the Optimization of a Discrete Loss

### v1

@@ 3.2 Diffusion equation
where we have set $\beta=${10}^{4}\text{\,}$\$. This inverse problem is ill-posed because the initial conditions are unknown and the diffusion process is inherently not time reversible [11]. Thus, many different initial states can produce indistinguishable noisy measurements at later times. The MAP associated with eq. 17 is shown on fig. 3, and we can observe deviations between the inferred field at $t=0$ and the initial conditions used to generate the dataset. This deviation is expected at early times, as explained above.
@@ 3.3 Reaction-diffusion equation
The dataset $\mathcal{D}$ is synthetically generated for each case using eq. 20 with the numerical solution of eq. 19 at time $T$, with initial conditions described by eq. 22. The solution is obtained with $\mathbf{\theta}_{\mathrm{ref}}=(2L/3,L/3)$, $L=1$, $T=0.5$, and $\rho=8$. The diffusion coefficient takes values $D(x,y)\in\{0.005,0.1\}$. Finally, we use $n_{x}=n_{y}=64$ points along each spacial dimension and a time step $\Delta t=$6.1\text{\times}{10}^{-4}\text{\,}$\$.
@@
We use the Bayesian annealed sequential importance sampling (BASIS) algorithm [38] to sample the posterior distribution of $\mathbf{\theta}=(x_{0},y_{0})$, with parameters $\beta_{\text{BASIS}}=0.2$, $\gamma_{\text{BASIS}}=1$, $l_{max}=1$ and 256 samples. The posterior distribution of the initial position $\mathbf{\theta}=(x_{0},y_{0})$ given the data is computed for different values of $\sigma$ in both cases 1 and 2. Samples of these distributions are shown in fig. 5. A higher value of $\sigma$ means less trust in the data, hence a higher uncertainty in $\mathbf{\theta}$. Using samples from these posterior distributions, we can reconstruct the field $u_{f}$ at $t=T$ and evaluate their uncertainties by using
@@ 3.4 Reconstruction of tumor cells concentration in the brain
Medical imaging such as MRI can detect regions on high tumor concentration in the brain, known as gross tumor volume (GTV). However, regions with lower tumor cell density, often containing microscopic signatures of a disease, are not detected using standard imaging modalities. Accurately estimating the full tumor concentration field is essential for developing effective radiation therapy. Recently, several methods have been proposed to estimate the tumor concentration field. The newly introduced method of GliODIL [32] uses ODIL to estimate the tumor concentration satisfying a reaction-diffusion PDE while matching MRI data of brain tumors.
@@
with zero flux at the boundaries of the brain $\nabla u\cdot\mathbf{n}=0$ on $\partial\Omega$, $\mathbf{n}$ being the normal vector at the brain boundaries. Furthermore we assume initial conditions of the form
@@
where $0<\tau_{\text{lo}}<\tau_{\text{up}}<1$ are parameters. The solution $u$ is discretized on a uniform grid that spans a space 50% larger than that spanned by the tumor at time $T$ with $64\times 64\times 64$ grid points in space and 128 points in time. Equation 26 is discretized using the Crank-Nicolson scheme, as explained in ref. [32] and in appendix A.
@@
As in the previous section, we use the BASIS algorithm to create 512 samples from the posterior distribution $P(x_{0},y_{0},z_{0}|\mathcal{D})$, with $\beta_{\text{BASIS}}=0.2$, $\gamma_{\text{BASIS}}=1.5$, $l_{max}=1$. For each sample, the optimization is performed over 5000 epochs with a learning rate of ${10}^{-3}\text{\,}$, reduced by half if no progress was made for 50 consecutive iterations, with a minimal value of ${10}^{-4}\text{\,}$. Each sample is mapped on a H100 GPU and takes about 15 minutes to evaluate. We have used MRI scans that are part of the dataset used in ref. [32].
@@ Acknowledgments
The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. We thank Sebastian Kaltenbach (Harvard University) and Michal Balcerak (University of Zürich) for their valuable input about Bayesian inference and tumor growth model.
@@ Appendix A Discretization used in GliODIL
Equation 26 is discretized on a uniform grid with field values $u_{i,j,k}^{n}$, where $n$ are time indices and $i$, $j$, and $k$ are spacial indices. The diffusion and reaction terms are defined as

### v2

@@ 3.3 Diffusion equation
where we have set $\beta=${10}^{4}\text{\,}$\$. We compute the MAP by maximizing eq. 20 using the LBFGS optimizer [36] as implemented in PyTorch. Each call to the optimizer performs up to 20 internal quasi-Newton iterations with a strong Wolfe line search. We run 500 such optimization steps, retaining a curvature history of size 100 throughout. The initial step size is set to $5\text{\times}{10}^{-4}\text{\,}$. Computing the Laplace approximation for this case takes about $10\text{\,}\mathrm{s}$ on a consumer-grade laptop. This inverse problem is ill-posed because the initial conditions are unknown and the diffusion process is inherently not time reversible [25]. Thus, many different initial states can produce indistinguishable noisy measurements at later times. The MAP associated with eq. 19 is shown on fig. 5, and we can observe deviations between the inferred field at $t=0$ and the initial conditions used to generate the dataset. This deviation is expected at early times, as explained above.
@@ 3.4 Reaction-diffusion equation
The dataset $\mathcal{D}$ is synthetically generated for each case using eq. 22 with the numerical solution of eq. 21 at time $T$, with initial conditions described by eq. 24. The solution is obtained with $\mathbf{\theta}_{\mathrm{ref}}=(2L/3,L/3)$, $L=1$, $T=0.5$, and $\rho=8$. The diffusion coefficient takes values $D(x,y)\in\{0.005,0.1\}$. Finally, we use $n_{x}=n_{y}=64$ points along each spatial dimension and a time step $\Delta t=$6.1\text{\times}{10}^{-4}\text{\,}$\$.
@@
We use the Bayesian annealed sequential importance sampling (BASIS) algorithm [56] to sample the posterior distribution of $\mathbf{\theta}=(x_{0},y_{0})$, with parameters $\beta_{\text{BASIS}}=0.2$, $\gamma_{\text{BASIS}}=1$, $l_{max}=1$ and 256 samples. Each sample is computed with 20000 epochs of the Adam optimizer with learning rate ${10}^{-3}\text{\,}$ and required approximately $8\text{\,}\mathrm{min}$ on a consumer-grade laptop and $2.3\text{\,}\mathrm{min}$ on a A100 GPU. The posterior distribution of the initial position $\mathbf{\theta}=(x_{0},y_{0})$ given the data is computed for different values of $\sigma$ in both cases 1 and 2. Samples of these distributions are shown in fig. 7. A higher value of $\sigma$ means less trust in the data, hence a higher uncertainty in $\mathbf{\theta}$. Using samples from these posterior distributions, we can reconstruct the field $u_{f}$ at $t=T$ and evaluate their uncertainties by using
@@ 3.5 Reconstruction of tumor cells concentration in the brain
Medical imaging such as MRI can detect regions on high tumor concentration in the brain, known as gross tumor volume (GTV). However, regions with lower tumor cell density, often containing microscopic signatures of a disease, are not detected using standard imaging modalities. Accurately estimating the full tumor concentration field is essential for developing effective radiation therapy. Recently, a large body of work has focused on biophysical tumor growth models and their use for patient-specific inference from medical imaging. These approaches typically rely on reaction-diffusion or anisotropic diffusion models, sometimes coupled with mass effect, and have been used for tasks such as tumor growth prediction, parameter estimation, and image-based personalization (see e.g., Clatz et al. [13], Hogea et al. [22, 23], Swanson et al. [51], Biegler et al. [8], Lipková et al. [35], Menze et al. [38], Miniere et al. [39], LaMonica et al. [32] and references therein). The newly introduced method of GliODIL [7] uses ODIL to estimate the tumor concentration satisfying a reaction-diffusion PDE while matching MRI data of brain tumors.
@@
with zero flux at the boundaries of the brain $\nabla u\cdot\mathbf{n}=0$ on $\partial\Omega$, $\mathbf{n}$ being the normal vector at the brain boundaries. The no-flux condition is enforced by setting the diffusion coefficient $D$ to zero outside brain tissue, so that diffusive fluxes vanish at the interface between brain matter and surrounding regions. Furthermore we assume initial conditions of the form
@@
where $0<\tau_{\text{lo}}<\tau_{\text{up}}<1$ are parameters. The solution $u$ is discretized on a uniform grid that spans a space 50% larger than that spanned by the tumor at time $T$ with $64\times 64\times 64$ grid points in space and 128 points in time. Equation 28 is discretized using the Crank-Nicolson scheme, as explained in Balcerak et al. [7] and in appendix A.
@@
As in the previous section, we use the BASIS algorithm to create 512 samples from the posterior distribution $P(x_{0},y_{0},z_{0}|\mathcal{D})$, with $\beta_{\text{BASIS}}=0.2$, $\gamma_{\text{BASIS}}=1.5$, $l_{max}=1$. For each sample, the optimization is performed over 5000 epochs with the Adam optimizer and a learning rate of ${10}^{-3}\text{\,}$, reduced by half if no progress was made for 50 consecutive iterations, with a minimal value of ${10}^{-4}\text{\,}$. Each sample is mapped on a H100 GPU and takes about 15 minutes to evaluate. We have used MRI scans that are part of the dataset used in Balcerak et al. [7].
@@ Acknowledgments
The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. We thank Sebastian Kaltenbach (Harvard University) and Michal Balcerak (University of Zürich) for their valuable input about Bayesian inference and tumor growth model. This work was supported by the Harvard Data Science Initiative (HDSI) and Amazon Web Services (AWS) under Award Number A59479.
@@ Appendix A Discretization used in GliODIL
Equation 28 is discretized on a uniform grid with field values $u_{i,j,k}^{n}$, where $n$ are time indices and $i$, $j$, and $k$ are spatial indices. The diffusion and reaction terms are defined as